ROTATIONAL SPECTRUM OF CIS-CIS HOONO

\underline{BRIAN J. DROUIN}, CHARLES E. MILLER, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, CA 91109; JULIANE L. FRY, Department of Chemistry, California Institute of Technology, Pasadena, CA 91125.

The pure rotational spectrum of cis-cis peroxynitrous acid, HOONO, has been observed for the first time. Over 220 transitions, sampling states up to J′ = 67 and K′_{a} = 31, have been fitted with an RMS uncertainty of 48.4 kHz. The experimentally determined rotational constants agree well with ab initio values for the cis-cis conformer, a five-membered ring formed by intramolecular hydrogen bonding. The small, positive inertial defect, Δ = 0.075667(60) amu Å^{2}, and lack of any observable torsional splittings in the spectrum indicate that cis-cis HOONO exists in a well-defined planar structure at room temperature. Analysis of the rotational spectrum of the deuterated isotopologue, DOONO, confirms the planar molecular structure and yields substitution coordinates for the hydrogen atom.